OpenStudy (kkutie7):

I don't think that I' doing this problem right so can someone please help/ check my work?

OpenStudy (kkutie7):

question: for any complex number z, show that: \[a)Re(z)=\frac{(z+z^{*})}{2}\] \[b)Im(z)=\frac{(z-z^{*})}{2i}\]

OpenStudy (kkutie7):

ok so I guess I make up two simple complex numbers. I'm confused why the function are labeled as Re(real) and Im(imaginary)

OpenStudy (kkutie7):

so I guess: \[z=1+i\] \[z^{*}=1-i\]

OpenStudy (kkutie7):

\[a)Re(z)=\frac{((1+i)+(1-i))}{2}\] \[\frac{2}{2}=1\]

OpenStudy (kkutie7):

\[b)Im(z)=\frac{((1+i)-(1-i))}{2i}\] \[\frac{1+i-1+i}{2i} \rightarrow \frac{2i}{2i}=1\]

OpenStudy (andrewyates):

You're correct, except you're proving it for the specific case where \(z = 1+i\). Do the exact same thing, except with \(z = a+bi\).

OpenStudy (andrewyates):

\(Re(z)\) is the real part of \(z\). For example, the real part of \(5+2i\) is \(5\). \(Im(z)\) is the imaginary part of \(z\). The imaginary part of \(5+2i\) is 2.

OpenStudy (kkutie7):

like this? \[a)Re(z)=\frac{((a+bi)+(a-bi))}{2}=\frac{2a}{2}=a\] \[b)Im(z)=\frac{((a+bi)-(a-bi))}{2i}=\frac{2bi}{2i}=b\]

OpenStudy (andrewyates):

Yep, exactly.

OpenStudy (kkutie7):

so when im asked, for example, give a geometric interpretation of why z x z* is a real number i can use the same concept?

OpenStudy (kkutie7):

or should i be plotting stuff

OpenStudy (andrewyates):

I'm not exactly sure how you would give a geometric interpretation of why it's true, but yes, you can use the same concept to show why it's real.

OpenStudy (andrewyates):

Yes, they probably want you plotting the numbers if they say geometric. The problem is \(z \ \cdot \ z*\) is multiplication, so you can't use vector addition. Are you familiar with the polar form of complex numbers? \(z=Re^{i\theta}\)

OpenStudy (kkutie7):

euler's rule yes

OpenStudy (kkutie7):

so what have two different z and turn them into their polar forms then multiply them together?

OpenStudy (andrewyates):

Yes, so \(z=Re^{i\theta}\) and \(z^* = Re^{-i\theta}\). You can prove the latter by using Euler's formula.

OpenStudy (andrewyates):

This way, you can show that multiplying by a complex number with angle \(\theta\) is the same as rotating by \(\theta\) in the complex plane. That's your geometric interpretation.

OpenStudy (kkutie7):

so by using what you have: \[Re^{i\theta}*Re^{-i\theta} \rightarrow Re^{0} \rightarrow R\]

OpenStudy (andrewyates):

Yes, except it's \(R^2\)

OpenStudy (kkutie7):

oh right and R^{2} is a real number

OpenStudy (andrewyates):

Yeah, and you can show that geometrically by plotting them with their angles labeled and then show how multiplying by \(z^*\) just "rotates \(z\) backwards" onto the real axis.

OpenStudy (kkutie7):

this is were i'm confused

OpenStudy (andrewyates):

Okay, so when you're plotting a complex number in polar form (\(z=Re^{i\theta}\)), you are basically plotting a vector with magnitude R at an angle \(\theta\) with the x-axis. Does this make sense?

OpenStudy (kkutie7):

i under stand how to plot them but what do you mean by it moving backwards

OpenStudy (andrewyates):

When you multiplied in polar form, you added \(-\theta\) to \(\theta\) in the exponent: \[Re^{i\theta}∗Re^{−i\theta} = R^2e^{i(\theta - \theta)}\] The new complex number is "rotated backwards" by theta. Do you see this in the exponent?

OpenStudy (kkutie7):

so the angle is just being subtracted by theta. i guess im getting confused because i can really see it on a graph. i just imagine that theta=> zero after subtraction so it should be on the x axis, but when i think of plotting re^(-itheta) i see it as the i value along negative y and theta being huge

OpenStudy (kkutie7):

would you mind helping my draw this so i understand

OpenStudy (kkutie7):

my thought